# Regularity of Einstein $5$-manifolds via $4$-dimensional gap theorems

### v1

@@ 5.5. Geodesics on limits of Einstein $5$-manifolds
We conclude by classifying all geodesics in the limit space $X$.
In a seminal work [CoN12], Colding and Naber proved the Hölder continuity of tangent cones along any limit geodesic that arises as the limit of geodesics in the sequence. More recently Deng [Den25] generalized the estimate to any geodesic in $X$, that is a curve locally minimizing the distance $d$.
@@
It suffices to consider the case where the entire $\sigma((0,l))\in\mathcal{S}$. By the stratification Theorem 1.15, there must be some $\Gamma$ and $t_{0}\in(0,l)$ such that $\sigma(t_{0})\in\mathcal{S}^{1}_{\Gamma}$ since $\mathcal{S}^{0}$ is countable. Then by Theorem 5.13 and isolation Theorem 1.9, there exists some $\theta_{0}>0$ such that the tangent cone at $\sigma(t)$ is isometric to $\mathds{R}\times\mathds{R}^{4}/\Gamma$ for any $t\in(t_{0}-\theta_{0},t_{0}+\theta_{0})$. By connectedness this implies that any interior point has the unique tangent cone $\mathds{R}\times\mathds{R}^{4}/\Gamma$. Therefore we can use Theorem 1.16 to prove that the interior of this geodesic is removable in the sense of Theorem 1.16. ∎
Theorem 1.19 therefore provides a complete classification of the interior of all geodesics. We say that a geodesic is of $\Gamma$-type if all the tangent cones are $\mathds{R}\times\mathds{R}^{4}/\Gamma$ (we allow $\Gamma=\{id\}$). We now turn to the behavior of the end points. We prove that the tangent cones at end points can only have worse singularities. Similar results can be formulated for any curves thanks to Theorem 1.16.
@@
Let $\sigma:[0,l]\to X$ be a geodesic of $\Gamma$-type in $X$. Let $C(Z)$ be a tangent cone at $\sigma(0)$. Then $Z$ has at least one singularity of the type $\mathds{R}^{4}/\Gamma^{\prime}$ with $|\Gamma^{\prime}|\geq|\Gamma|$.

### v2

@@ 5.5. Geodesics on limits of Einstein $5$-manifolds
We conclude by classifying all geodesics in the limit space $X$. We first recall the Hölder continuity estimate along geodesic in [CoN12] and also [Den25].
@@
It suffices to consider the case where the entire $\sigma((0,l))\in\mathcal{S}$. By the stratification Theorem 1.15, there must be some $\Gamma$ and $t_{0}\in(0,l)$ such that $\sigma(t_{0})\in\mathcal{S}^{1}_{\Gamma}$ since $\mathcal{S}^{0}$ is only countable. Then by Theorem 5.13 and isolation Theorem 1.9, there exists some $\theta_{0}>0$ such that the tangent cone at $\sigma(t)$ is isometric to $\mathds{R}\times\mathds{R}^{4}/\Gamma$ for any $t\in(t_{0}-\theta_{0},t_{0}+\theta_{0})$. By connectedness this implies that any interior point has the unique tangent cone $\mathds{R}\times\mathds{R}^{4}/\Gamma$. Therefore we can use Theorem 1.16 to prove that the interior of this geodesic is removable in the sense of Theorem 1.16. ∎
Theorem 1.21 therefore provides a complete classification of the interior of all geodesics. We say that a geodesic is of $\Gamma$-type if the tangent cones at all interior points are $\mathds{R}\times\mathds{R}^{4}/\Gamma$ (we allow $\Gamma=\{id\}$). We now turn to the behavior of the end points. We prove that the tangent cones at end points can only have worse singularities. Similar results can be formulated for any curves thanks to Theorem 1.16.
@@
Let $\sigma:[0,l)\to X$ be a geodesic of $\Gamma$-type in $X$. Let $C(Z)$ be a tangent cone at $\sigma(0)$. Then $Z$ has at least one singularity of the type $\mathds{R}^{4}/\Gamma^{\prime}$ with $|\Gamma^{\prime}|\geq|\Gamma|$.